The following equation is used to predict quarterly demand: Yt = 350 - 2.5 t, where t = 0 is the second quarter of last year. Quarterly indices are: Q1 = 1.4; Q2 = 0.8; Q3 = 1.1; and Q4 = 0.7. What is the forecast for the fourth quarter of this year? a. 335.0 b. None of the answers provided c. 201.0 d. 234.5 e. 268.0
Answer #1
Qtr t Forcast Indices Forcast Using indices
350-2.5t
a b c=350-(2.5*b) d e=c*d)
1 3 342.5 1.4 479.5
2 4 340 0.8 272
3 5 337.5 1.1 371.25
4 6 335 0.7 234.5
Correct Option: b
Let me know your queries. Please rate.